OpenStudy (anonymous):

AC Method y^2-6y-40 O.o

OpenStudy (anonymous):

y^2-6y-40=0 (y-10)(y+4)=0 (y-10)=0 and (y+4)=0 solve for y... :D

OpenStudy (anonymous):

To use the AC method, find AC that multiplies to -40. (ac = 1 * -40). Two numbers that multiply out to -40 and add up to -6 (b) is the factor form.

OpenStudy (anonymous):

Thanks :)